OpenStudy (anonymous):

Hi, I'm having trouble simplifying this equation and am stuck after solving the inner tan parts: (2tan(pi/12))/(1-tan(pi/12))

OpenStudy (anonymous):

this is how far I got: \[(2(2-\sqrt(3)))/(\sqrt(3)-1)\]

OpenStudy (mertsj):

Better find an identity that is equal to: \[\frac{2tanx}{1-tanx}\]

OpenStudy (mertsj):

Are you sure it is not 1-tan^2(x) on the bottom?

OpenStudy (anonymous):

yes I'm sure mertsj, it's 1-tan(pi/12) in the denominator

OpenStudy (anonymous):

and I can't think of any identity that would match that, I've looked at my identity's page

OpenStudy (anonymous):

I typed it up a little more prettier: \[2\tan(\pi/12)/1-\tan(\pi/12)\]

OpenStudy (anonymous):

Wolfram Alpha says it should reduce to \[\sqrt{3} -1\]

OpenStudy (mertsj):

Did you try going back to sin and cos?

OpenStudy (anonymous):

No, not yet mertsj, I was thinking that'd be too cumbersome

OpenStudy (anonymous):

I'd be especially confused on what to do in the denominator then, 1 - sin/cos(pi/12) ? Do both the sin and cos take pi/12 then or what?

OpenStudy (mertsj):

2sin(pi/12)]cos(pi/12)-sin(pi/12)

OpenStudy (anonymous):

bunch of half angle formula right?

OpenStudy (anonymous):

mertsj: is that bracket after the first sin(pi/12) supposed to be a division sign?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

satellite: I'm not exactly sure, just some simplification problems with trig functions

OpenStudy (anonymous):

what is wrong with your first answer? looks ok to me

OpenStudy (anonymous):

I don't think it's simple enough, I am wondering how to go from that to \[\sqrt{3} -1 \]

OpenStudy (anonymous):

\[\frac{2(2-\sqrt{3})}{\sqrt{3}-1}\] maybe rationalize the denominator by multiplying by \[\sqrt{3}+1\] top and bottom

OpenStudy (anonymous):

your denominator will be 2, and that will cancel with the 2 in the numerator so last job is \[(2-\sqrt{3})(\sqrt{3}+1)\]

OpenStudy (anonymous):

yes that is where I am. Do you mean to multiply by \[\sqrt{3} - 1 instead of \sqrt{3} + 1 \] ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but how will the one with the plus cancel out the \sqrt(3) - 1 in the denominator ?

OpenStudy (anonymous):

if you want to rationalize the denominator, multiply by \(\sqrt{3}+1\) because then yo will get \[(\sqrt{3}-1)(\sqrt{3}+1)=3-1=2\]

OpenStudy (anonymous):

as \((a-b)(a+b)=a^2-b^2\)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

and as i said the 2 in the denominator will cancel with the 2 in the numerator, then multiply \[(2-\sqrt{3})(\sqrt{3}+1)\] and i bet you get your answer

OpenStudy (anonymous):

Everytime I get \[\sqrt{3} + 1\] instead of the \[\sqrt{3}-1\] answer that I got from wolfram alpha

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

yes, I tried it multiiple times

OpenStudy (anonymous):

minus square root of 3 times square root of 3 is -3 and 2 times 1 is 2 and 2 - 3 = -1

OpenStudy (anonymous):

\[(2-\sqrt{3})(\sqrt{3}+1)\] \[2\sqrt{3}+2-\sqrt{3}\sqrt{3}-\sqrt{3}\] \[2\sqrt{3}+2-3-\sqrt{3}\] \[\sqrt{3}-1\]

OpenStudy (anonymous):

Ah, I didn't completely finish distributing the first F in FOIL that is why I missed the 2 from 2 minus sqrt 3 times sqrt 3. Thank you so much